40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 1014     Entry time: Wed Oct 1 02:54:03 2008
Author: rob 
Type: Update 
Category: Locking 
Subject: bad 

Tried the spring-y side tonight with a discouraging lack of progress. There were several locks of DRMI+2ARMs with
the +f2 (springy) sideband resonating in the DRM, but they weren't very stable. Moving to just the DRMI and resonating
the +f2, in order to tune up the acquisition and the handoff to the double demod signals, revealed the problem that the
DRM just won't stay locked on the +f2 sideband. It locks quickly, but only for a few seconds. This is different from the
behaviour with the -f2 sideband, which locks quickly and stably. In theory, the two sidebands should behave similarly.
It could be problems with HOMs in the recycling cavities, and so we may try changing the modulation frequency slightly.
ELOG V3.1.3-